if cuberoot(a)+cuberoot(b)+cuberoot(c)=0 then what is the value of (a+b+c)^3= ??? 
 

Grade:

1 Answers

ADARSH SINGH
5 years ago
SINCE A3+ B3+C3 = 0.
 
WE KNOW THE IDENTITY THAT ,
 
(A+B+C)3 = A3+B3+C3 +3(A+B)(B+C)(C+A)
 
SO IN THIS CASE , 
 
(A+B+C)3= 3(A+B)(B+C)(C+A)
 
HOPE YOU GOT YOUR ANSWER.
THANK YOU .
HAVE A NICE DAY .
AND PLEASE REPLY FOR ANY FURTHER QUERIES.